OpenStudy (anonymous):

Anne takes 2 days to paint a house. Joe takes 4. if they do it together how long would it take them ?

OpenStudy (anonymous):

if u suppose that the two guys do the work in x hours 1 part of the work done by Anne = 1/2 1 part of the work done by joe = 1/4 1 part of the work done by both = 1/x 1/2 +1/4 = 1/x 3/4 = 1/x x= 4/3

OpenStudy (anonymous):

yees thank you so much!!!!!!

OpenStudy (amistre64):

i just wonder howmany times this house needs to be painted.... whats the lead count on it by now?

OpenStudy (anonymous):

:D very true :P